OpenStudy (anonymous):

Roger attends a local carnival, he spends $8.75 on food and $1.25 for each ride, he uses $15.00 all together. A.Write an equation to represent this situation B.How many rides did Rogar ride?

OpenStudy (anonymous):

A) We know that the total amount of money is $15: \[t=15\] We also know that he spent $8.75 on food: \[f=8.75\] and $1.25 per ride: \[r=1.25n\] So the total must be \[15=8.75+1.25n\] B) Solve for n \[15=8.75+1.25n\] \[6.25=1.25n\]\[n=5\] He rode 5 rides.

OpenStudy (texaschic101):

nicely done :)

OpenStudy (anonymous):

Thanks for the response <3